# A Survey on Human-Centered Evaluation of Explainable AI Methods in Clinical Decision Support Systems

### v3

@@ 3.1 CDSS adopting Intrinsic Interpretable Models
Intrinsic interpretable models are easily testable in the medical field for CDSS due to their transparency, as no post-hoc evaluations are necessarily required to perform model evaluations. Examples in this scope are the works from Brennan et al. brennan2019comparing in Critical Care medicine, and Gombolay et al. gombolay2024effects in Neurology.  brennan2019comparing validated the usability of MyRiskSurgery, an analytical CDSS based on generalized linear models and random forests to estimate postoperative risk complications bihorac2019mysurgeryrisk. Through think-aloud studies, a set of physicians evaluated risk complications for a set of 8-10 individual cases in two case studies such as diagnosing acute kidney injury and Intensive Care Unit (ICU) admission, with the system displaying the top three contributing features for each prediction. As a result, physicians experienced notable improvements in their risk assessment, which led to a substantial net improvement in reclassification rates from a no XAI evaluation (+12% and +16%, respectively), finding the system to be beneficial.
Finally, gombolay2024effects conducted a study to evaluate the impact of various XAI techniques on clinician interactions with CSDSS in pediatric and adult neurological disorder. Clinicians were exposed to different XAI methodologies, including decision trees, crowd-sourced agreement, case-based reasoning, probability scores, counterfactual reasoning, feature importance, templated language, and no explanations. The study found that decision trees were perceived as the most explainable. Interestingly, increasing neurology experience and perceived explainability were associated with degraded performance, raising the concern that there is not one-size-fits-all XAI medical approach, but strong CDSS personalization is warranted.
@@
With only two studies surveyed, MyRiskSurgery showed positive perception and seamless usability by clinicians brennan2019comparing. While,  gombolay2024effects raised the concern that XAI customizations are required for better understandability based on the specific applied case.
@@ 3.2 CDSS adopting Post Hoc XAI Methodologies
In the field of Critical Care medicine,  barda2020qualitative developed a CDSS to prevent in-hospital mortality in Pediatric Intensive Care Units (PICU). The CDSS leveraged a random forest and SHAP values on top. In general, SHAP methods quantify each feature’s contribution to a model’s prediction by computing Shapley values from cooperative game theory (for an exhaustive review please refer to lundberg2017shap; shapley1953value). Overall, feedback from focus group sessions with clinicians indicated that the proposed solution was perceived as useful. However, the capacity to digest the information varied across clinical roles and their level of predictive modeling knowledge, with preferred solutions that required less cognitive effort. Next,  he2024vms developed the Visualization for Model Sensemaking and Selection (VMS), an interactive tool designed to aid in the selection of machine learning models to predict patients’ Length of Stay (LOS) in ICU. The system employs a combination of decision tree, random forest, LSTM, and GRU models, with SHAP values used to provide both local and global explanations of feature importance. Feedback from a user study study involving 16 medical professionals indicated that the tool enhanced CDSS interpretability, although it was noted that there was a learning curve associated with its use.
@@
With the same goal, and adopting a similar SHAP methodology,  jung2025evaluating found that clinicians aligned more closely with the CDSS after receiving explanations, though full alignment was not reached. Next,  hur2025comparison developed the Precision Maximum Surgical Blood Ordering Schedule-Thoracic Surgery (pMSBOS-TS) to predict personalized red blood cell demand in thoracic surgery patients. pMSBOS-TS employs a backbone using boosting methods, integrating SHAP to highlight the top three contributing features, and was evaluated with 63 clinicians, including surgeons and physicians, who participated in a repeated-measures experiment using clinical vignettes to evaluate three explanation formats: Results Only (RO), Results with SHAP visualizations (RS), and Results with SHAP visualizations plus clinical explanations (RSC). The RSC format significantly improved trust, satisfaction, and acceptance compared to the other formats.
@@
Works discussed so far have primarily focused on perioperative or intraoperative critical care. As for postoperative care,  abraham2023integrating developed a ML-augmented tool designed to predict complications such as acute kidney injury, delirium, pneumonia, among others. The tool employs SHAP values to provide interpretations of risk factors, and was evaluated with 17 clinicians, including anesthesiologists, surgeons, and critical care physicians, who participated in cognitive walkthroughs and interviews to evaluate the tool’s usability and integration into clinical workflows. The findings indicated a high level of agreement between ML-generated risk predictions and clinicians’ manual risk assessments. Also, clinicians appreciated the tool’s ability to streamline report preparation and enhance care planning by highlighting patient-specific risks.
@@
As for sepsis treatment,  sivaraman2023ignore examined clinicians’ interactions with an AI-based interpretable sepsis treatment recommender. The system was modeled using XGBoost, and SHAP as feature explainer. Through a think-aloud study involving 24 intensive care clinicians, the analysis revealed that explanations improved clinicians’ perceptions of the AI’s usefulness. In particular, the study identified four distinct behavior patterns among clinicians summarised as: Ignore, Negotiate, Consider, and Rely. Specifically, in the Negotiate group, clinicians displayed a tendency to weigh and prioritize aspects of the AI recommendations, suggesting that treatment decisions in the ICU may involve partial reliance on AI recommendations, with potential implications for the effectiveness of chosen treatments. Overall, clinicians who found the AI recommendations most useful saw it as additional evidence alongside their assessments. Next, in neurocritical care, Bienefeld2023 presented the development and evaluation of the Delayed Cerebral Ischemia Predictor (DCIP), a CDSS designed to forecast delayed cerebral ischemia in patients with aneurysmal subarachnoid hemorrhage. The system combined Extremely Randomized Trees and SHAP values to display static and dynamic risk contributors, and was evaluated with 112 clinicians and developers using a mixture of surveys, focus groups, and think-aloud studies. Results showed that while developers emphasized model interpretability, clinicians sought clinical plausibility, leading to design tensions around the utility of SHAP explanations.
@@
All the works in critical care reviewed so far adopted SHAP as explainability framework. On the contrary, kumarakulasinghe2020evaluating adopted LIME for sepsis prediction. In general, LIME methods approximate a model’s behavior around a specific instance by fitting an interpretable surrogate model to locally perturbed samples ribeiro2016should. Developers trained several models and applied LIME to the best-performing random forest. Physicians evaluated agreement with sepsis predictions and independently ranked key features before revealing LIME explanations. Afterward, they rated satisfaction on a five-star Likert scale. Physicians agreed with the model in 87% of cases, with LIME aligning on at least two features for 69%. Overall, 78% were satisfied and 68% showed positive attitudes toward explainability, though inconsistencies in LIME limited trust.
@@
Next, in the field of Endocrinologykovalchuk2022three proposed a three-staged CDSS for diabetic diagnosis as follows: (1) Basic Reasoning, i.e., implementing healthcare practices by referencing existing policies, (2) Data-Driven, i.e., applying machine learning models in the context of the previous stage, and (3) Explanation, in which SHAP values were applied on top of (2) to provide better recommendations. Unexpectedly, the basic reasoning case (1) showed higher understandability, followed by the explanation case (2), and the data-driven case (3). These findings suggest that adding explanations is useful, but well-grounded, known conventional clinical practices may still outperform AI-driven solutions. Similarly,  bhattacharya2023directive developed a visualization-based CDSS to monitor and evaluate the risk of diabetes onset. The system was built upon a logit model and incorporated multiple explanation types, including rule-based data-centric explanations, SHAP, and counterfactual explanations generated using DiCE mothilal2020dice. A preliminary focus group study revealed that while SHAP-based explanations were informative, practitioners found them too complex and instead preferred simpler, data-centric visualizations such as bar charts combined with comparative patient data and actionable suggestions. Similarly, counterfactual explanations were valued for their ability to present example-based recommendations, supporting actionable insights. In contrast, feature-importance explanations (SHAP) were seen as less actionable. Finally,  chari2023informing developed a CDSS to enhance the interpretability of risk predictions for chronic kidney disease (CKD) among type-2 diabetes mellitus (T2DM) patients. The system employs LLMs, specifically BERT and SciBERT, to generate contextual information that align AI predictions with clinical contexts, and integrates SHAP values for deeper understanding of the risk factors involved. The evaluation involved an expert panel of 4 medical professionals who assessed the system’s explanations for 20 prototypical patients, highlighting the need to provide actionable insights within clinical workflows.
@@
Across diverse medical domains, we documented a similar trend of how post hoc XAI methods were used. For instance, in Oncologygu2020acase proposed a CDSS for breast cancer recurrence prediction that combined XGBoost with SHAP-based visualizations within a case-based reasoning framework. A survey of 32 oncologists revealed a favorable reception, highlighting SHAP’s usefulness compared to alternative interpretability methods. A similar emphasis on SHAP was seen from  jing2025development that, in the field of Geriatrics, developed an interpretable XGBoost model for frailty risk prediction using real-time health data and validated scales. Through a mixed-methods evaluation involving healthcare professionals and data scientists, the study reported strong interpretability (mean rating 4.30/5) and high user satisfaction. In contrast,  matthiesen2021clinician, in Cardiology, developed a CDSS employing a random forest model visualized via LIME to predict ventricular fibrillation risk, finding that while the explanations increased clinicians’ confidence, they did not alter clinical decisions. Finally, in Gastroenterologyrajashekar2024human integrated LLMs with traditional ML by developing GutGPT, combining natural language reasoning with ICE and ALE plots to generate interpretable narratives for gastrointestinal bleeding risk prediction. Broadly speaking, ICE plots visualize how a model prediction for individual instances changes as a single feature varies. More globally, ALE plots summarize the average effect of a feature on model predictions by integrating local changes. Users perceived such explanations positively, suggesting the promise of hybrid approaches that blend human-understandable text with quantitative interpretability.
@@
Across the 15 studies reviewed, eight in Critical Care, three in Endocrinology, and the remaining in Oncology, Cardiology, Geriatrics, and Gastroenterology, perceptions of XAI-based CDSS were predominantly positive. Twelve studies employed SHAP, and two used LIME for patient-specific assessments of feature contributions. This functionality, particularly in time-sensitive domains such as critical care, appeared valuable in augmenting clinicians’ confidence and perceived reliability of AI-driven recommendations. Among the studies reporting positive perceptions, explanations were generally valued for their clarity, ease of interpretation, and perceived cognitive support in clinical reasoning barda2020qualitative; kumarakulasinghe2020evaluating; gu2020acase; jing2025development; rajashekar2024human.
Several works emphasized that despite the favorable attitude, certain limitations persisted, including the need for user adaptation, indicating a learning curve associated with use he2024vms, as well as for solutions demanding less cognitive effort barda2020qualitative. Finally, neutral perceptions typically reflected a cautious attitude toward integration, where explanations were appreciated as complementary evidence rather than decisive inputs sivaraman2023ignore; matthiesen2021clinician; kovalchuk2022three. While interpretability was recognized as beneficial, traditional heuristics and clinical experience often retained precedence. Moreover, studies reported that explanation formats varied in their perceived actionability, in which counterfactual and data-driven ones being seen as more actionable than feature-importance visualizations such as SHAP bhattacharya2023directive. In aggregate, the evidence suggests a consistently positive or neutral reception of XAI in clinical decision support, with no reported negative perceptions.
@@ 3.3 CDSS adopting XAI Methodologies in Deep Learning
In these fields, ellenrieder2023promoting extended the work of pumplun2023bringing that leveraged a CNN and the integrated gradients framework to produce heatmaps of image regions to highlight cancer malignancy probabilities. In technical terms, the integrated gradients method attributes a model’s prediction to its input features by accumulating the gradients of the output with respect to the input from a baseline to the actual input, producing a fine-grained, axiomatic feature attribution map. In collaboration with radiologists, a think-aloud study evaluated learning outcomes across four UI designs: non-explainable high-performing (NEHP), explainable high-performing (EHP), non-explainable low-performing (NELP), and explainable low-performing (ELP). The results confirmed four hypotheses: (H1) NEHP yielded modest learning gains; (H2) EHP produced greater gains; (H3) NELP induced false learning; and (H4) ELP mitigated false learning, showing that explainability may significantly learning.
@@
Next, using CAM methods,  cabitza2025fromoracular proposed a Judicial AI, a framework focusing on diagnosing vertebral fractures using X-ray images, employing CAM activation maps as primary explainability technique. In general, CAM methods identify the spatial regions within an input image that most contribute to a model’s prediction by linearly combining convolutional feature maps weighted by class-specific coefficients, typically obtained from architectures employing global average pooling. These maps highlight anatomical regions supporting different diagnostic option. The system was evaluated through a survey involving 16 medical professionals, including spine surgeons and radiologists, who assessed 18 complex X-ray cases. Results indicated that Judicial AI improved diagnostic accuracy and clinician confidence, particularly among experienced users, while less experienced users found the system more cognitively demanding. Similarly, famiglini2024evidence conducted a study to evaluate the impact of CAM in support to a CDSS using a CNN to detect thoracolumbar fractures from X-rays. Their evaluation through an online questionnaire administered to 16 medical professionals, including spine surgeons and musculoskeletal radiologists, indicated that traditional coloring schemes in CAM consistently yielded higher diagnostic accuracy and increased physician confidence.
@@
Next, GradCAM methods have also been used for similar tasks. Broadly speaking, Grad-CAM methods generalize CAM by using the gradients of the target output with respect to feature maps to derive these weights, allowing class-discriminative localization in convolutional architectures without requiring architectural modifications. In this subfield,  rainey2024operationalizing conducted a study to evaluate the impact of a CDSS to analyze musculoskeletal radiographs. The AI behind provided two types of feedback: binary diagnoses and visual explanations using GradCAM. The study involved 12 reporting radiographers, who were presented with 18 complete examinations. The evaluation focused on the participants’ trust in the AI system and their propensity to change decisions based on AI feedback, indicating a significant correlation between agreement with AI feedback and trust, particularly with the GradCAM visualizations. However, participants showed a preference for simple binary feedback over heatmaps, as the latter often led to disagreement.
@@
Always in radiology,  singla2023medical proposed a Generative Adversarial Network (GAN)-based CDSS for X-ray images dissection with multiple explanation frameworks. Broadly speaking, GANs are a class of machine learning models in which two neural networks, the generator and the discriminator, are trained simultaneously in a competitive framework to produce increasingly realistic synthetic data. The CDSS was evaluated through a human-centered study involving 12 diagnostic radiology residents, who assessed the understandability, decision justification, visual quality, identity preservation, and overall helpfulness of the explanations. The results indicated that the counterfactual explanations significantly improved the participants’ understanding of the classifier’s decisions compared to other methods like saliency maps, suggesting that visual explanatory formats are not always better off than more standard approaches. Next, with a more comprehensive evaluation of XAI techniques,  ihongbe2024evaluating developed a CDSS based on two CNNs to diagnose pneumonia from chest X-rays and CT scans, employing GradCAM and LIME to provide explanations. A human-centered evaluation was conducted through surveys involving medical professionals, assessing the XAI tools in terms of clinical relevance, revealing a generally positive perception of XAI systems among participants, although there was a noted lack of awareness regarding their practical applications. Notably, GradCAM was preferred over LIME for its superior performance.
@@
Next, in Oncological Imagingsabol2020explainable adopted a Cumulative Fuzzy Class Membership Criterion (CFCMC) to predict colorectal cancer. The CDSS was based on two segmentation systems for whole-slide histopathological images: one using a standalone CNN and the other using the CFCMC classifier on CNN-extracted features to provide explanations. Through surveys with 14 pathologists, the CFMC showed higher acceptability compared to a plain CNN, pointing out that explainable classifiers improve AI usability and reliability in medical settings. Finally, in Dental Radiology,  kayadibi2025ai proposed an innovative Explainable Mandibular Third Molar Convolutional Neural Network (E-mTMCNN) architecture to enhance the detection of mandibular third molars in panoramic radiography, aiming to improve treatment in dentistry. E-mTMCNN utilized a CNN model, equipped with LIME to provide localized, interpretable insights into the model’s predictions. Ten specialized surveyed dentists assessed the system, indicating high clinical acceptance and confidence in the system’s reliability.

### v4

@@ 3.1 CDSS adopting Intrinsic Interpretable Models
Intrinsic interpretable models are easily testable in the medical field for CDSS due to their transparency, as no post-hoc evaluations are necessarily required to perform model evaluations. Examples in this scope are the works from Brennan et al. Brennan et al. (2019) in Critical Care medicine, and Gombolay et al. Gombolay et al. (2024) in Neurology.  Brennan et al. (2019) validated the usability of MyRiskSurgery, an analytical CDSS based on generalized linear models and random forests to estimate postoperative risk complications Bihorac et al. (2019). Through think-aloud studies, a set of physicians evaluated risk complications for a set of 8-10 individual cases in two case studies such as diagnosing acute kidney injury and Intensive Care Unit (ICU) admission, with the system displaying the top three contributing features for each prediction. As a result, physicians experienced notable improvements in their risk assessment, which led to a substantial net improvement in reclassification rates from a no XAI evaluation (+12% and +16%, respectively), finding the system to be beneficial.
Finally, Gombolay et al. (2024) conducted a study to evaluate the impact of various XAI techniques on clinician interactions with CSDSS in pediatric and adult neurological disorder. Clinicians were exposed to different XAI methodologies, including decision trees, crowd-sourced agreement, case-based reasoning, probability scores, counterfactual reasoning, feature importance, templated language, and no explanations. The study found that decision trees were perceived as the most explainable. Interestingly, increasing neurology experience and perceived explainability were associated with degraded performance, raising the concern that there is not one-size-fits-all XAI medical approach, but strong CDSS personalization is warranted.
@@
With only two studies surveyed, MyRiskSurgery showed positive perception and seamless usability by clinicians Brennan et al. (2019). While,  Gombolay et al. (2024) raised the concern that XAI customizations are required for better understandability based on the specific applied case.
@@ 3.2 CDSS adopting Post Hoc XAI Methodologies
In the field of Critical Care medicine,  Barda et al. (2020) developed a CDSS to prevent in-hospital mortality in Pediatric Intensive Care Units (PICU). The CDSS leveraged a random forest and SHAP values on top. In general, SHAP methods quantify each feature’s contribution to a model’s prediction by computing Shapley values from cooperative game theory (for an exhaustive review please refer to Lundberg and Lee (2017); Shapley (1953)). Overall, feedback from focus group sessions with clinicians indicated that the proposed solution was perceived as useful. However, the capacity to digest the information varied across clinical roles and their level of predictive modeling knowledge, with preferred solutions that required less cognitive effort. Next,  He et al. (2024) developed the Visualization for Model Sensemaking and Selection (VMS), an interactive tool designed to aid in the selection of machine learning models to predict patients’ Length of Stay (LOS) in ICU. The system employs a combination of decision tree, random forest, LSTM, and GRU models, with SHAP values used to provide both local and global explanations of feature importance. Feedback from a user study study involving 16 medical professionals indicated that the tool enhanced CDSS interpretability, although it was noted that there was a learning curve associated with its use.
@@
With the same goal, and adopting a similar SHAP methodology,  Jung et al. (2025) found that clinicians aligned more closely with the CDSS after receiving explanations, though full alignment was not reached. Next,  Hur et al. (2025) developed the Precision Maximum Surgical Blood Ordering Schedule-Thoracic Surgery (pMSBOS-TS) to predict personalized red blood cell demand in thoracic surgery patients. pMSBOS-TS employs a backbone using boosting methods, integrating SHAP to highlight the top three contributing features, and was evaluated with 63 clinicians, including surgeons and physicians, who participated in a repeated-measures experiment using clinical vignettes to evaluate three explanation formats: Results Only (RO), Results with SHAP visualizations (RS), and Results with SHAP visualizations plus clinical explanations (RSC). The RSC format significantly improved trust, satisfaction, and acceptance compared to the other formats.
@@
Works discussed so far have primarily focused on perioperative or intraoperative critical care. As for postoperative care,  Abraham et al. (2023) developed a ML-augmented tool designed to predict complications such as acute kidney injury, delirium, pneumonia, among others. The tool employs SHAP values to provide interpretations of risk factors, and was evaluated with 17 clinicians, including anesthesiologists, surgeons, and critical care physicians, who participated in cognitive walkthroughs and interviews to evaluate the tool’s usability and integration into clinical workflows. The findings indicated a high level of agreement between ML-generated risk predictions and clinicians’ manual risk assessments. Also, clinicians appreciated the tool’s ability to streamline report preparation and enhance care planning by highlighting patient-specific risks.
@@
As for sepsis treatment,  Sivaraman et al. (2023) examined clinicians’ interactions with an AI-based interpretable sepsis treatment recommender. The system was modeled using XGBoost, and SHAP as feature explainer. Through a think-aloud study involving 24 intensive care clinicians, the analysis revealed that explanations improved clinicians’ perceptions of the AI’s usefulness. In particular, the study identified four distinct behavior patterns among clinicians summarised as: Ignore, Negotiate, Consider, and Rely. Specifically, in the Negotiate group, clinicians displayed a tendency to weigh and prioritize aspects of the AI recommendations, suggesting that treatment decisions in the ICU may involve partial reliance on AI recommendations, with potential implications for the effectiveness of chosen treatments. Overall, clinicians who found the AI recommendations most useful saw it as additional evidence alongside their assessments. Next, in neurocritical care, Bienefeld et al. (2023) presented the development and evaluation of the Delayed Cerebral Ischemia Predictor (DCIP), a CDSS designed to forecast delayed cerebral ischemia in patients with aneurysmal subarachnoid hemorrhage. The system combined Extremely Randomized Trees and SHAP values to display static and dynamic risk contributors, and was evaluated with 112 clinicians and developers using a mixture of surveys, focus groups, and think-aloud studies. Results showed that while developers emphasized model interpretability, clinicians sought clinical plausibility, leading to design tensions around the utility of SHAP explanations.
@@
All the works in critical care reviewed so far adopted SHAP as explainability framework. On the contrary, Kumar. et al. (2020) adopted LIME for sepsis prediction. In general, LIME methods approximate a model’s behavior around a specific instance by fitting an interpretable surrogate model to locally perturbed samples Ribeiro et al. (2016). Developers trained several models and applied LIME to the best-performing random forest. Physicians evaluated agreement with sepsis predictions and independently ranked key features before revealing LIME explanations. Afterward, they rated satisfaction on a five-star Likert scale. Physicians agreed with the model in 87% of cases, with LIME aligning on at least two features for 69%. Overall, 78% were satisfied and 68% showed positive attitudes toward explainability, though inconsistencies in LIME limited trust.
@@
Next, in the field of EndocrinologyKovalchuk et al. (2022) proposed a three-staged CDSS for diabetic diagnosis as follows: (1) Basic Reasoning, i.e., implementing healthcare practices by referencing existing policies, (2) Data-Driven, i.e., applying machine learning models in the context of the previous stage, and (3) Explanation, in which SHAP values were applied on top of (2) to provide better recommendations. Unexpectedly, the basic reasoning case (1) showed higher understandability, followed by the explanation case (2), and the data-driven case (3). These findings suggest that adding explanations is useful, but well-grounded, known conventional clinical practices may still outperform AI-driven solutions. Similarly,  Bhattacharya et al. (2023) developed a visualization-based CDSS to monitor and evaluate the risk of diabetes onset. The system was built upon a logit model and incorporated multiple explanation types, including rule-based data-centric explanations, SHAP, and counterfactual explanations generated using DiCE Mothilal et al. (2020). A preliminary focus group study revealed that while SHAP-based explanations were informative, practitioners found them too complex and instead preferred simpler, data-centric visualizations such as bar charts combined with comparative patient data and actionable suggestions. Similarly, counterfactual explanations were valued for their ability to present example-based recommendations, supporting actionable insights. In contrast, feature-importance explanations (SHAP) were seen as less actionable. Finally,  Chari et al. (2023) developed a CDSS to enhance the interpretability of risk predictions for chronic kidney disease (CKD) among type-2 diabetes mellitus (T2DM) patients. The system employs LLMs, specifically BERT and SciBERT, to generate contextual information that align AI predictions with clinical contexts, and integrates SHAP values for deeper understanding of the risk factors involved. The evaluation involved an expert panel of 4 medical professionals who assessed the system’s explanations for 20 prototypical patients, highlighting the need to provide actionable insights within clinical workflows.
@@
Across diverse medical domains, we documented a similar trend of how post hoc XAI methods were used. For instance, in OncologyGu et al. (2020) proposed a CDSS for breast cancer recurrence prediction that combined XGBoost with SHAP-based visualizations within a case-based reasoning framework. A survey of 32 oncologists revealed a favorable reception, highlighting SHAP’s usefulness compared to alternative interpretability methods. A similar emphasis on SHAP was seen from  Jing et al. (2025) that, in the field of Geriatrics, developed an interpretable XGBoost model for frailty risk prediction using real-time health data and validated scales. Through a mixed-methods evaluation involving healthcare professionals and data scientists, the study reported strong interpretability (mean rating 4.30/5) and high user satisfaction. In contrast,  Matthiesen et al. (2021), in Cardiology, developed a CDSS employing a random forest model visualized via LIME to predict ventricular fibrillation risk, finding that while the explanations increased clinicians’ confidence, they did not alter clinical decisions. Finally, in GastroenterologyRajashekar et al. (2024) integrated LLMs with traditional ML by developing GutGPT, combining natural language reasoning with ICE and ALE plots to generate interpretable narratives for gastrointestinal bleeding risk prediction. Broadly speaking, ICE plots visualize how a model prediction for individual instances changes as a single feature varies. More globally, ALE plots summarize the average effect of a feature on model predictions by integrating local changes. Users perceived such explanations positively, suggesting the promise of hybrid approaches that blend human-understandable text with quantitative interpretability.
@@
Across the 15 studies reviewed, eight in Critical Care, three in Endocrinology, and the remaining in Oncology, Cardiology, Geriatrics, and Gastroenterology, perceptions of XAI-based CDSS were predominantly positive. Twelve studies employed SHAP, and two used LIME for patient-specific assessments of feature contributions. This functionality, particularly in time-sensitive domains such as critical care, appeared valuable in augmenting clinicians’ confidence and perceived reliability of AI-driven recommendations. Among the studies reporting positive perceptions, explanations were generally valued for their clarity, ease of interpretation, and perceived cognitive support in clinical reasoning Barda et al. (2020); Kumar. et al. (2020); Gu et al. (2020); Jing et al. (2025); Rajashekar et al. (2024).
Several works emphasized that despite the favorable attitude, certain limitations persisted, including the need for user adaptation, indicating a learning curve associated with use He et al. (2024), as well as for solutions demanding less cognitive effort Barda et al. (2020). Finally, neutral perceptions typically reflected a cautious attitude toward integration, where explanations were appreciated as complementary evidence rather than decisive inputs Sivaraman et al. (2023); Matthiesen et al. (2021); Kovalchuk et al. (2022). While interpretability was recognized as beneficial, traditional heuristics and clinical experience often retained precedence. Moreover, studies reported that explanation formats varied in their perceived actionability, in which counterfactual and data-driven ones being seen as more actionable than feature-importance visualizations such as SHAP Bhattacharya et al. (2023). In aggregate, the evidence suggests a consistently positive or neutral reception of XAI in clinical decision support, with no reported negative perceptions.
@@ 3.3 CDSS adopting XAI Methodologies in Deep Learning
In these fields, Ellenrieder et al. (2023) extended the work of Pumplun et al. (2023) that leveraged a CNN and the integrated gradients framework to produce heatmaps of image regions to highlight cancer malignancy probabilities. In technical terms, the integrated gradients method attributes a model’s prediction to its input features by accumulating the gradients of the output with respect to the input from a baseline to the actual input, producing a fine-grained, axiomatic feature attribution map. In collaboration with radiologists, a think-aloud study evaluated learning outcomes across four UI designs: non-explainable high-performing (NEHP), explainable high-performing (EHP), non-explainable low-performing (NELP), and explainable low-performing (ELP). The results confirmed four hypotheses: (H1) NEHP yielded modest learning gains; (H2) EHP produced greater gains; (H3) NELP induced false learning; and (H4) ELP mitigated false learning, showing that explainability may significantly learning.
@@
Next, using CAM methods,  Cabitza et al. (2025) proposed a Judicial AI, a framework focusing on diagnosing vertebral fractures using X-ray images, employing CAM activation maps as primary explainability technique. In general, CAM methods identify the spatial regions within an input image that most contribute to a model’s prediction by linearly combining convolutional feature maps weighted by class-specific coefficients, typically obtained from architectures employing global average pooling. These maps highlight anatomical regions supporting different diagnostic option. The system was evaluated through a survey involving 16 medical professionals, including spine surgeons and radiologists, who assessed 18 complex X-ray cases. Results indicated that Judicial AI improved diagnostic accuracy and clinician confidence, particularly among experienced users, while less experienced users found the system more cognitively demanding. Similarly, Famiglini et al. (2024) conducted a study to evaluate the impact of CAM in support to a CDSS using a CNN to detect thoracolumbar fractures from X-rays. Their evaluation through an online questionnaire administered to 16 medical professionals, including spine surgeons and musculoskeletal radiologists, indicated that traditional coloring schemes in CAM consistently yielded higher diagnostic accuracy and increased physician confidence.
@@
Next, GradCAM methods have also been used for similar tasks. Broadly speaking, Grad-CAM methods generalize CAM by using the gradients of the target output with respect to feature maps to derive these weights, allowing class-discriminative localization in convolutional architectures without requiring architectural modifications. In this subfield,  Rainey et al. (2024) conducted a study to evaluate the impact of a CDSS to analyze musculoskeletal radiographs. The AI behind provided two types of feedback: binary diagnoses and visual explanations using GradCAM. The study involved 12 reporting radiographers, who were presented with 18 complete examinations. The evaluation focused on the participants’ trust in the AI system and their propensity to change decisions based on AI feedback, indicating a significant correlation between agreement with AI feedback and trust, particularly with the GradCAM visualizations. However, participants showed a preference for simple binary feedback over heatmaps, as the latter often led to disagreement.
@@
Always in radiology,  Singla et al. (2023) proposed a Generative Adversarial Network (GAN)-based CDSS for X-ray images dissection with multiple explanation frameworks. Broadly speaking, GANs are a class of machine learning models in which two neural networks, the generator and the discriminator, are trained simultaneously in a competitive framework to produce increasingly realistic synthetic data. The CDSS was evaluated through a human-centered study involving 12 diagnostic radiology residents, who assessed the understandability, decision justification, visual quality, identity preservation, and overall helpfulness of the explanations. The results indicated that the counterfactual explanations significantly improved the participants’ understanding of the classifier’s decisions compared to other methods like saliency maps, suggesting that visual explanatory formats are not always better off than more standard approaches. Next, with a more comprehensive evaluation of XAI techniques,  E. Ihongbe et al. (2024) developed a CDSS based on two CNNs to diagnose pneumonia from chest X-rays and CT scans, employing GradCAM and LIME to provide explanations. A human-centered evaluation was conducted through surveys involving medical professionals, assessing the XAI tools in terms of clinical relevance, revealing a generally positive perception of XAI systems among participants, although there was a noted lack of awareness regarding their practical applications. Notably, GradCAM was preferred over LIME for its superior performance.
@@
Next, in Oncological ImagingSabol et al. (2020) adopted a Cumulative Fuzzy Class Membership Criterion (CFCMC) to predict colorectal cancer. The CDSS was based on two segmentation systems for whole-slide histopathological images: one using a standalone CNN and the other using the CFCMC classifier on CNN-extracted features to provide explanations. Through surveys with 14 pathologists, the CFMC showed higher acceptability compared to a plain CNN, pointing out that explainable classifiers improve AI usability and reliability in medical settings. Finally, in Dental Radiology,  Kayadibi et al. (2025) proposed an innovative Explainable Mandibular Third Molar Convolutional Neural Network (E-mTMCNN) architecture to enhance the detection of mandibular third molars in panoramic radiography, aiming to improve treatment in dentistry. E-mTMCNN utilized a CNN model, equipped with LIME to provide localized, interpretable insights into the model’s predictions. Ten specialized surveyed dentists assessed the system, indicating high clinical acceptance and confidence in the system’s reliability.